OpenStudy (anonymous):

0.3 x-0.2y=4 0.2x+0.3y=5/23 solve by using the elimination method

OpenStudy (anonymous):

multiply the first equation with 3 and the second equation with 2 and add them up. you will eliminate y

OpenStudy (anonymous):

IM STILL CONFUSED

OpenStudy (anonymous):

did you try doing what I said?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

what did you get? post it here

OpenStudy (anonymous):

I DONT GET IT SORRY

OpenStudy (anonymous):

well, you just said you tried doing what I said. I am asking you to post what you got so that we can see where you are getting stuck.

OpenStudy (anonymous):

oh. sorry, can't do your problems for you. ask someone else.

OpenStudy (anonymous):

0,3x-0,2y =4 ....(1) 0,2x+0,3y=5/23...(2) first multiple (1) with 20 so, being 6x-4y=80.... (1)' second multiple (2) with 30 so, 6x+3y=150/23....(2)' third eliminate (1)' with (2)' being -7y =80-150/23 -7y = 73,48 y = - 10,49 and then substitute y to (1)' so, 6x-4(-10,49) =80 6x+41,96=80 6x = 80-41,96 6x = 38,04 x =6,34

OpenStudy (anonymous):

SO HOW WOULD THE ANSWER APPEAR ( 6,34)